OpenStudy (anonymous):

evaluate the integral $$\int_R\int xy dx dy$$ over the region R covered by $$\frac{x^2}{a^2}+\frac{y^2}{b^2}= 1$$ in the first quadrant

OpenStudy (irishboy123):

\[\int\limits_{x = 0}^{a}\int\limits_{y = 0}^{ +b \sqrt { \left (1 - \frac{x^2}{a^2} \right )}} xy \ dy \ dx\]

OpenStudy (irishboy123):

you can param it too, but this is a pretty easy integral

OpenStudy (anonymous):

Thank you!!...struggled to find limits. :)

OpenStudy (irishboy123):

and this works too... \[\int\limits_{y =0}^{b} \int\limits_{x = 0}^{+a \sqrt{ 1 - \frac{y^2}{b^2} }} x \ y \ dx \ dy\]